OpenStudy (anonymous):

I need help, please help an object starts from rest at time t=0 and moves in the x direction with constant acceleration. The object travels 12m from time t=1 to time t=2. what is the acceleration of the object?

OpenStudy (anonymous):

Is the answer 24ms^-2

OpenStudy (anonymous):

Since the object starts from rest, you can express it's initial velocity in the period between t=1 and t=2 in terms of the acceleration: \[v(t) = v_0 + a*\Delta t\] \[v(1) = 0 +a*1 = a\] and substitute this result to the x(t) position equation: \[x(\Delta t) = v_0*\Delta t + 0.5*a*\Delta t^2 \] \[x(1->2) = a*1 + 0.5*a*1 = 1.5*a = 12\] and from here you get: \[a = 8 (m/\sec^2)\]

OpenStudy (anonymous):

Thank you Elb27 for your help